# https://neu-vi.github.io/HOI-Diff/

### v2

@@ 4.2 Comparisons with Existing Methods
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
User Study $\uparrow$
Real 0.096 0.324 7.244 - - -
@@
MDM* [45] 3.169 0.206 7.606 0.448 0.184 0.87%
PriorMDM* [41] 4.352 0.215 8.186 0.416 0.270 16.53%
InterDiff [54] 7.192 0.170 7.686 0.506 0.218 0.00%
Ours (Full) 1.501 0.235 7.391 0.347 0.182 82.60%
Table 2: Quantitative results on OMOMO dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
Real 1.539 0.193 4.833 - -
MDM* 9.155 0.137 5.353 0.768 0.191
PriorMDM 11.192 0.128 7.404 0.523 0.344
InterDiff 10.567 0.110 6.493 0.906 0.239
Ours 8.661 0.168 4.780 0.326 0.141
Figure 5: Qualitative comparisons of our approach and baseline methods on BEHAVE dataset. The bottom row, showcasing our method, demonstrates the generation of realistic 3D HOIs with plausible contacts, particularly evident in columns 2 and 4. This contrasts with the baselines, which fail to achieve a similar level of realism and contact plausibility in the interactions. As an additional visual aid, the mesh color gradually darkens over time to represent progression. (Best viewed in color.)
Quantitative Results. Sec. 4.2 reports the quantitative results on BEHAVE dataset [3]. Compared with the baseline methods, our full method achieves the best performance. Specifically, it achieves state-of-the-art results in both FID, R-precision, and Diversity, underscoring its ability to generate high-quality human motions in the context of coherently interacting with objects. The best Contact Distance also suggests that our approach can generate physically plausible HOIs, capturing the intricate interplay interactions between humans and objects.
We also show the results of our User Study in Tab. 4.2 to illustrate user preferences between our method and other baselines. In our user study, we sample 24 sequences from each method and 23 participants are asked to choose their most preferred generation results from these samples. This user study is designed to directly compare user preferences across different methods. The result indicates a strong preference for our method: it is favored over the baselines in a substantial majority of 82.60% of the cases.
Tab. 4.2 presents the quantitative results on the OMOMO dataset. Our method consistently outperforms other baselines by a considerable margin across all metrics. Notably, due to the distinctiveness of objects in the training and testing sets, the results in Tab. 4.2 indicate the effectiveness of our approach in generalizing to unseen objects, proving superior performance compared to other models.
Table 3: Ablation studies of our model’s variants on the BEHAVE dataset. The right arrow normal-$\rightarrow$ means closer to real data is better.
FID $\downarrow$
R-precision $\uparrow$
Diversity $\rightarrow$ Contact Distance $\downarrow$ Foot Skate Ratio $\downarrow$
User Study $\uparrow$
w/o Interaction Correction
Ours w/o CM 2.825 0.201 7.850 0.524 0.265 0.00%
Ours w/o pretrain 2.414 0.226 7.915 0.402 0.158 3.90%
@@
Ours 1.808 0.217 7.718 0.416 0.205 11.69%
w/ Interaction Correction
Ours w/o ${M}^{o}$ CM 3.206 0.205 7.909 0.365 0.310 0.00%
Ours ${}^{Joint}$ 3.412 0.195 7.878 0.421 0.342 -
Ours w/o $G_{con}$ 1.983 0.205 7.826 0.417 0.196 -
Ours w/o $G_{sta}$ 1.636 0.213 7.785 0.367 0.181 -
Ours w/o $G_{smo}$ 1.665 0.205 7.840 0.370 0.182 -
Ours (Full) 1.501 0.235 7.391 0.347 0.182 84.41%
@@
Figure 6: Visual results of different variants of our model in ablation studies on the BEHAVE dataset [3].
Qualitative Results. We showcase qualitative comparisons, rendered with SMPL [27] shapes, between our approach and the baseline methods in Fig. 5. It is observed that the generated HOI motion by other baselines lacks smoothness and realism, where the object may float in the air (e.g., the toolbox in Fig. 5 (b)). Furthermore, these baseline methods struggle to accurately capture the spatial relationships between humans and objects (e.g., the chair in Fig. 5 (e)). In stark contrast, our approach excels in creating visually appealing and realistic HOIs. Notably, it adeptly reflects the intricate details outlined in text descriptions, capturing both the nature of the interactive actions and the specific body parts involved (e.g., raising the trash bin with the right hand in Fig. 5 (a)).

### v3

@@ Appendix D Additional Details of Evaluation Metrics
Model DBDM APDM
Parameters (·106) 8.82 38.92
Figure 8: Model Parameters. The number of learned parameters of our two core architectures.
@@
Method MDM* PriorMDM* Ours (Full)
Time (s) 32.3 38.6 118.0
Component APDM DBDM Interaction Correction
Time (s) 24.2 46.4 47.4
Figure 9: Inference Time (on NVIDIA A5000 GPU). We report the inference time for baselines, our full method, and its key components.
Appendix E Inference Time
In Table 9, we provide the inference times for both baselines and our full method, including its key components. All measurements were conducted using an NVIDIA A5000 GPU. Training an additional model for affordance information and using classifier guidance for interaction correction do contribute to increased inference costs. However, despite the longer inference time, our complete method notably enhances the accuracy of 3D HOIs generation.
Params (M) FID $\downarrow$ R-precision (Top-3) $\uparrow$
MDM 49.85 6.98 0.36
Ours (Full) 47.74 1.62 0.46
Table 4: With comparable model size, the performance results of MDM and Ours (Full).
Figure 10: Effect of different control strengths for classifier guidance. (a) We use equal strengths of $\tau_{1}=1,\tau_{2}=1$ to perturb the predicted mean of human motion and object motion, respectively. (b) We use different strengths of $\tau_{1}=1,\tau_{2}=100$ for the perturbation. We can see that different strengths work better.
Appendix F Additional Ablation Studies
Different perturbing times in classifier guidance. As discussed in Sec. 3.4, in the later stage of classifier guidance, diffusion models tend to strongly attenuate the introduced signals. Therefore, we iteratively perturb the predicted mean of motion for $K$ times at the final denoising step. In Figure 7, we present the ablation results, illustrating the impact of different numbers of perturbations. Notably, we observe that employing 100 perturbations leads to re-convergence and yields the desired results.
Different guidance strength. As detailed in Sec. 3.4, we employ distinct control strengths for classifier guidance, considering the varying feature densities in predicted human and object motion. Rather than employing equal control strengths, we opt to assign a higher control strength to object motion, allowing it to closely align with human contact joints, as illustrated in Figure 10.
Different model with comparable model size. Although our method involves a slightly larger number of model parameters, our model is specifically designed for HOI generation. As seen in the  Table 4, if we attempt to scale MDM* to the same model size, its performance remains subpar.
Appendix G User Study
For each method, we select 15 prompts from the BEHAVE dataset and 10 prompts from the OMOMO dataset, covering various interaction types and object items. We sample twice with each prompt to gather a total of 50 results. 40 participants are asked to choose their most preferred generation results from these samples. This user study requires pairwise comparisons of our method with other baseline on generated interaction quality, as shown in Figure 17. The results in Figure 12 indicate strong preference for our method: it is favored over the baselines in 89.6% (Ours vs. MDM*), 73.8% (Ours vs. PriorMDM*) and 95.3% (Ours vs. Interdiff).
@@ Appendix G User Study
Figure 11: Perceptual User Study. Most participants prefer our method over the baselines.
[…]
@@
Figure 12: Visual results of different variants of our model in ablation studies.
Appendix H Additional Qualitative Results
In this section, we present additional qualitative results showcasing the model’s performance evaluated on the OMOMO dataset, and the effectiveness of APDM.
Figure 13: Additional qualitative evaluation on OMOMO dataset. Given object geometry and text description, our method can generate high-quality human-object interactions even for the unseen objects (tripod, smalltable, suitcase).
Figure 14: Additional qualitative evaluation on unseen objects.
Figure 15: Visual results of estimated contact points. Our APDM, trained on the BEHAVE dataset, can accurately estimating contact positions for objects based on textual descriptions. Furthermore, it showcases the capability to generalize to unseen objects in the OMOMO dataset, as demonstrated in the last row.
Qualitative results on OMOMO dataset. We present additional qualitative results on the OMOMO dataset, rendered with SMPL [SMPL:2015] shapes, as shown in Figure 13. It is evident that our method can generalizes effectively to unseen objects and produce realistic 3D human-object interactions.
Qualitative results of APDM. To verify the accuracy of estimated contact points on object surface, we provide additional visual results in Figure 15. It can be seen that our method can predict realistic and practical contact points based on text descriptions. With APDM, we even can generate different interactions with the same object based on the input description, as shown in the Figure 16.
Generalization capability. To verify the model’s generalization capability, except of unseen object test on OMOMO dataset, we also downloaded several objects from Sketchfab [^3] https://sketchfab.com/, adjusted them to a reasonable scale, and used them as inputs. As shown in Figure 14, our model successfully establishes reasonable HOI contact with these previously unseen objects.
Figure 16: Leveraging the power of the APDM module, our method can generate diverse HOIs for the same object using different contacting body parts and contact points.
Figure 17: An example question for our text-to-hoi user study.
Appendix I Annotation for BEHAVE Dataset
Text Annotating Process. Initially, we manually annotate the interaction types and the specific human body parts involved, delineating actions like “lift” associated with the “left hand” or “hold” involving “two hands”. Subsequently, to generate complete sentences, we leverage the capabilities of GPT-3.5 to assist in formulating the entirety of the description.
Table 5: Examples of our annotated textual descriptions for the BEHAVE dataset rephrased by GPT-3.5 [chatgpt].
Object Textual Descriptions
backpack A person is carrying the backpack in front.
The person is raising a backpack with his right hand.
The person at the front presently has control over the backpack.
chairwood A person is using the chairwood for sitting.
(wooden chair) The person is propelling the chairwood on the ground.
Someone is hoisting a chairwood by his left hand.
tablesquare A person is lifting the tablesquare, utilizing his left hand.
(square table) Someone is clutching onto a tablesquare from the front.
An individual is moving the tablesquare back and forth.
boxlong A person is gripping the boxlong from the front.
(long box) A person is raising the boxlong using his left hand.
Someone hoists the boxlong with his left hand.
toolbox Someone is grasping the toolbox upfront.
The person has a firm hold on the toolbox with his right hand.
A person is gripping the toolbox with his left hand.
yogaball A person is shifting a yogaball back and forth on the floor using his hands.
The person is occupying a yogaball.
A person is employing an yogaball to engage in an upper body game.
Examples of Annotated Textual Descriptions. In Table 5, we showcase a selection of our annotated textual descriptions for the BEHAVE dataset [bhatnagar22behave].
Figure 18: Analysis of word frequency We count the occurrences of each interaction verb from all text descriptions to illustrate their respective frequencies.
Analysis of Annotated Textual Descriptions. All text descriptions encompass 36 distinct interaction verbs associated with 20 different objects. Figure 18 illustrates the frequency of each verb, indicating their respective occurrences.
Affordance Data. Our affordance data includes 8-dimensional human contact labels and object contact points. We employ chamfer distance to measure the distance between all human body joints and object surface points. Following a predefined distance threshold $\gamma=0.12$, we identify the 8 contact points on the object surface corresponding to the 8 primary human body joints. Subsequently, we derive the human contact labels by encoding the indexes of contact joints into an 8-dimensional vector represented by binary values.
Appendix J Additional Details of OMOMO Dataset
The OMOMO dataset comprises data captured for a total of 15 objects. Adhering to their official split strategy depicted in [li2023object](Figure 5), we allocate 10 objects for training and 5 objects for testing. This split allows us to further evaluate the model’s generalization ability to new objects. Notably, the OMOMO dataset itself provides text annotation, and we use GPT-3.5 to add subjects to it and embellish it appropriately. For affordance data, we preprocess it the same way we handle BEHAVE.
Appendix K Common Questions
Why use Skeletal Pose Representation rather than SMPL parameters? Most state-of-the-art text-to-motion methods adopt the skeletal pose representation proposed by Guo_2022_CVPR, demonstrating excellent performance and stability. While some works [azadi2023make] argue that SMPL parameters [SMPL:2015] contains shape and global information, it does not generate as smooth motions as skeletal-based approaches. Consequently, we adopt the skeletal pose representation and aim to leverage strong pose priors from the pretrained text-to-motion model [tevet2023human] to ensure the authenticity of generated human motion.
Can we handle multi-phase interactions between humans and objects? Due to the lack of fine-grained textural descriptions in the current 3D HOI dataset, we primarily consider only one interaction phase. However, we have found that an LLM can still reason well for multiple phases given a template such as: You will be given a sentence that describes an interaction between a person and an object across multiple phases. Your task is to divide the interaction into phases based on the state of the object and determine the state for each phase. If the object is being moved by the person during a phase, output the number 0. If the object remains stationary during a phase, output the number 1.
For example, given the text description: The box is on the ground. A person is picking up the box and holding it forward, then putting the box towards the table. The box is on the table" The result from GPT-3.5-turbo: "Phase 1: The box is on the ground - State: 1 (stationary); Phase 2: The person is picking up the box and holding it forward - State: 0 (moved); Phase 3: The person is putting the box towards the table - State: 0 (moved); Phase 4: The box is on the table - State: 1 (stationary). We will address the generation of multiple phases of 3D HOI in future work.
Can we generate hand motion with articulated fingers? The BEHAVE and OMOMO datasets do not capture and provide raw hand parameters, despite utilizing SMPLH and SMPLX models to fit human body meshes for rendering. Consequently, in this paper, we focus solely on whole-body human motion, excluding articulated hand and finger movements.
Why do we use large language models (LLMs) to predict object state based on the input description? We aim to leverage LLMs for inferring object states, and our results demonstrate that they perform efficiently and effectively. As shown in the Table 6, we evaluated the performance of object state prediction with GPT-3.5-turbo [chatgpt] and obtained an average precision of 95.6% on the validation set, with an average response time of 0.518 seconds. The results suggest that GPT-3.5-turbo is sufficiently accurate without adding significant overhead. We also evaluated the prediction performance using other LLMs, including Gemini-1.5-Pro-Exp-0801 [reid2024gemini] (99.4%, 0.259s), Gemma-2-27B [team2024gemma] (98.6%, 0.522s), and LLaMA-2-13B [touvron2023llama] (94.4%, 0.521s), the latter two being publicly available.
Acc (%) $\uparrow$ Time (s) $\downarrow$
GPT-3.5 95.6 0.518
Gemini-1.5-Pro-Exp-0801 99.4 0.259
Gemma-2-27B 98.6 0.522
LLaMA-2-13B 99.4 0.259
APDM + MLP 79.5 2.420
Table 6: LLMs’ inference accuracy (Acc) and average inference time (Time) on object state prediction.
To further validate the effectiveness of the LLM module, we modified the APDM module by adding an MLP head to predict the object status. The newly added MLP takes in the features consisting of object geometry information and CLIP embeddings. We used an MSE loss. We got average precision 79.5% and average time 2.42s for this design on the validation set, which is significantly worse than the results of GPT-3.5-turbo (95.6%, 0.518s), Gemma-2-27b (98.6%, 0.522s), Gemini-1.5-Pro-Exp-0801 (99.4%, 0.259s) and LLaMA-2-13B (4.4%, 0.521s).
In future work, we believe the LLM can play a more important role in 3D HOI, e.g. providing high-level instruction for more complex human-object interactions, and our initial use of the LLM offers insights into its potential applications and how it can be effectively utilized.
Appendix L Supplementary Video
Beyond the qualitative results presented in the main paper, our supplementary materials offer comprehensive demos that provide an in-depth visualization of our task, further showcasing the effectiveness of our approach.
In these demonstrations, we highlight the better performance of our method, HOI-Diff, in producing diverse and realistic 3D HOIs while maintaining adherence to physical validity. Notably, the visualizations show that HOI-Diff consistently generates smooth, vivid interactions, accurately capturing human-object contacts.
Additionally, we present the visual ablation results and emphasize the significance and effectiveness of our affordance-guided interaction correction, underscoring its substantial impact on improving the overall performance and quality of the generated 3D HOIs.
Appendix M Limitations
The existing datasets for 3D HOIs are limited in terms of action and motion diversity, posing a challenge for synthesizing long-term interactions in our task. Furthermore, the effectiveness of our model’s interaction correction component is contingent on the precision of affordance estimation. Despite simplifying this task, achieving accurate affordance estimation remains a significant challenge, impacting the overall performance of our model. A promising direction for future research involves integrating a sophisticated affordance model pre-trained on an extensive 3D object dataset, along with text prompts. Such an advancement could significantly enhance the realism and accuracy of human-object contact in our model, leading to more natural and precise HOIs synthesis.
Appendix N Social Impacts
On the positive side, it may offers the research community valuable insights into understanding human behaviors. On the negative side, it remains uncertain whether individuals can be identified solely based on their poses and movements. However, compared to traditional input images of people, this method poses a lower risk of invading personal privacy.